### v1

Abstract
Ring systems have recently been discovered around several small bodies in the outer Solar System through stellar occultations. While such measurements provide key information on ring geometry and dynamical interactions, little is known about their composition, grain-size distribution, origin, lifetime, or evolutionary pathways. Here we report near-infrared observations from the James Webb Space Telescope (JWST) of a stellar occultation by the Centaur (10199) Chariklo, providing unprecedented constraints on the material properties of a small-body ring system and offering insights into their origin and evolution. These measurements reveal that Chariklo’s inner dense ring contains predominantly micrometre-sized particles and exhibits a significant increase in opacity compared to previous observations, suggesting active replenishment events. Most strikingly, the outer ring shows a much weaker near-infrared occultation signature than in earlier visible-light detections. This discrepancy may indicate ongoing material loss, implying that the outer ring is transient, or it may reflect wavelength-dependent opacity consistent with a dusty structure dominated by 0.2–0.5 m silicate grains. These scenarios, not mutually exclusive, point to an unprecedented level of complexity in small-body ring systems, unlike anything observed around other minor bodies in the Solar System.
keywords:
KBOs, TNOs, Centaurs, stellar occultations, JWST, Chariklo, ringsStellar occultations are a powerful tool in planetary astronomy, offering unique opportunities to probe the properties of distant solar system bodies and their environments with exceptional precision. When a celestial object, such as an asteroid, passes in front of a star, the brief dimming of starlight reveals information about the occulting body’s size, shape, and surroundings [1, 2]. This technique has also enabled the detection and characterisation of tenuous atmospheres [3, 4, 5], or has placed stringent upper limits on their presence [2]. Notably, stellar occultations have led to the discovery and study of ring systems not only around the giant planets [6], but more recently around small bodies, including the Centaurs Chariklo [7] and Chiron [8, 9, 10], the dwarf planet Haumea [11], and the trans-Neptunian object Quaoar [12, 13]. Among these, Chariklo was the first small Solar System body discovered to host a ring system, composed of two sharply confined components, C1R and C2R, orbiting at 390 and 405 km from its centre, respectively [7, 14, 15]. Yet, despite extensive occultation campaigns and multiple studies on ring dynamics [16, 17, 18], the composition, grain-size distribution and origin of the rings have remained poorly constrained and have defied a clear explanation for over a decade. Constraining the composition of rings around small Solar System bodies remains challenging, as their proximity and faintness preclude the isolation of ring spectra using current remote spectroscopy [19]. However, wavelength-dependent optical properties of the grains, which lead to different occultation opacities in the visible and near-infrared, may provide direct constraints on grain properties [20], as it has been proposed for Haumea’s ring system [21].
To investigate the fine structure and composition of Chariklo’s rings beyond the limitations of ground-based observations, we targeted a stellar occultation using the James Webb Space Telescope (JWST). Predicting such an event from a space-based observatory is inherently challenging (see Section 1.1 for details); nevertheless, we successfully predicted and observed a stellar occultation by Chariklo on 18 October 2022. Remarkably, the JWST chord passed just 7.4 km from the body’s centre (Extended Data Figures), allowing us to detect the ring system (see also Section 1.4). No additional dips in brightness were observed that would indicate the presence of other material or orbiting satellites. This observation was conducted simultaneously in two near-infrared bands, centred at 1.5 and 3.2 m (F150W2 and F322W2), including the first occultation detection ever beyond 3 m, which is impossible from the ground. Exactly the same JWST setup was used later in the observations of the occultation event by the ring system of Quaoar [22]. Although the radial locations of both Chariklo’s rings in 2022 remain consistent with earlier measurements, they also exhibit significant variations compared to previous observations. For clarity, we adopt the terms first contact ( contact) and second contact ( contact) to refer to the two ring-plane crossings, each comprising an ingress and an egress: Ingress1–Egress1 and Ingress2–Egress2, respectively. This terminology will be used throughout the manuscript.
Chariklo’s inner dense ring, C1R, is clearly detected in the JWST light curves, exhibiting pronounced diffraction spikes at ingress and egress (see Figure 1), indicative of sharp edges and strong confinement, morphological features reminiscent of Uranus’s rings [23, 24]. Interestingly, C1R is found to be more opaque than before: the pre-JWST measurements provide an average normalized opacity of for this ring, while considering only the JWST measurements gives , indicating a highly significant difference, corresponding to a z-score of z = 4.2 and probability of . In contrast, the outer ring, C2R, displays an unexpectedly weak signal in the JWST measurements. After applying an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction (see Section 1.4), C2R was marginally detected at 1 in the F150W2 data, but not in the F322W2 filter, despite both measurements being performed simultaneously and sampling the same ring segment. Compared to previous ground-based visible range observations, this suggests particularly low opacities for C2R.
To unravel the origin of these differing opacities observed across epochs and wavelengths, we explore three basic scenarios that could explain these phenomena: 1) Azimuthally variable rings, where JWST consistently sampled a sparser or denser segment compared to earlier measurements; 2) Changes in the ring material, involving an increase or decrease in the total amount of material and/or the effective cross-section of grains, but without wavelength dependent grain properties; 3) Grains with wavelength-dependent optical properties, where composition and grain size effects are responsible for the observed variations.
In the case of a fragmented ring, alternating low- and high-density segments produce azimuthal variability in the observed ring structure, affecting both opacity and equivalent width. While previous occultation measurements yield roughly consistent values for these parameters, the JWST observations reveal significantly different opacity and equivalent width values. Using a simple azimuthal ring density model (see Section 1.5), we estimate that the probability of sampling only low-density (C2R) or high-density (C1R) segments during both ingress and egress in the JWST observations is very low, with p0.004 for C1R and p0.002 for C2R.
Changes in the effective cross-section in occultation measurement can be best described by the equivalent width (), which combines opacity, radial width, and the sensitivity and time resolution of the instrument. is proportional to the amount of material responsible for the extinction within the occulted segment (see Section 1.4 for its formal definition). We carried out a dedicated study of the C2R ring, combining the 2022 JWST occultation measurements (e.g., opacities, equivalent widths) with all available ground-based data from 2013, 2014, and 2017 (see Extended Data Tables and Section 1.4 for details). While no clear trend is seen in between 2013 and 2017, a marked decrease is observed between 2017 and 2022 for C2R (Figure 2, Panel (b)), potentially revealing a decade-long temporal evolution. This apparent fading of the ring may indicate that C2R is actively dispersing. Such depletion could result from collisions among particles, producing smaller grains that are more easily lost. Without effective confinement, these grains are subject to increasing orbital eccentricities due to radiation pressure, leading their orbits to intersect the inner ring [25]. Several known dusty rings in the Solar System exhibit temporal variability in opacity, morphology, or confinement, such as Saturn’s D, E, and F rings, Neptune’s Adams arcs, and Uranus’s ring. For instance, Saturn’s D ring has shown measurable radial shrinkage and brightness changes on decadal timescales, likely driven by micrometeoroid bombardment or electromagnetic forces [26]. The E ring, sourced by Enceladus’s plumes, varies with the moon’s activity and solar cycle [27]. Neptune’s Adams arcs display pronounced morphological evolution and changing brightness over months to years, pointing to a dynamically unstable configuration [28]. Uranus’s ring also shows opacity changes possibly linked to variable confinement mechanisms [29]. These examples illustrate that dusty, low-mass rings can undergo rapid and significant evolution, reinforcing the interpretation of Chariklo’s outer ring as a transient, evolving structure. Strikingly, the JWST opacity and equivalent width of C1R shows an opposite trend, a considerable increase (see Panel (a) in Figure 2, Extended Data Tables, and also [14, 15]). However, the change of in C1R () is approximately 10 times greater than that in C2R (), leaving further questions on the origin of the increase. In C1R, this could be due either to an overall increase in the amount of material or to the rise in the effective cross-section of grains, for example, through collisions between larger particles or rocks inside the ring, producing a cascade of smaller grains.
Since JWST probes a wavelength range where stellar occultations have rarely been observed, the discrepancy with visible-light data may also reflect the optical properties of the ring particles—specifically, a dominance of sub-micron grains whose scattering efficiency decreases at longer wavelengths. Even in the pre-JWST observations, a possible sign of wavelength dependence is the significantly different opacities of the 2013 and 2017 occultation measurements of C2R, using the Danish Z (1 m) and Danish Visual (0.55 m) bands (see Figure 3). To assess whether these variations reflect a genuine wavelength dependence, we adopt a radiative transfer approach using various grain types (e.g., water ice, olivine, carbon) and particle sizes, and compare the resulting extinction values with all available C2R occultation data (see Extended Data Tables and Extended Data Tables) to constrain plausible compositions consistent with the observations. A detailed description of this modelling can be found in Section 1.7. We find that for C2R, a wavelength dependence driven by different optical properties of the ring particles cannot be ruled out. Models dominated by silicates over water reproduce the detection in the F150W2 filter and the non-detection in the F322W2 filter, with pyroxene-rich compositions fitting both the visible and near-infrared opacities when grain sizes lie between 0.2 and 0.5 m (Figure 3). However, we emphasise that these compositional inferences depend on the assumed grain-size distribution and should not be viewed as uniquely limiting the ring’s material composition. In any case, if C2R is indeed dominated by micron-sized dust, its population may be analogous to that of Uranus’s ring, which consists predominantly of similar grains [30]. The ring likely originates from collisions involving a moonlet or a belt of larger parent bodies, potentially metres to kilometres in size, that generate fine particulate debris [31]. Similar dusty ring structures are seen co-orbiting with small Saturnian moons such as Janus, Anthe, Methone, and Pallene, where dust is believed to originate either from the satellites themselves or from co-orbiting larger bodies [32]. A comparable dusty ring structure has also been proposed around Makemake [25].
To explore the wavelength-dependent opacity scenario for C1R, we used a similar radiative transfer calculation, composed of a mixture of two materials with different compositions and potentially different grain sizes (see Section 1.7 and Figure 5). The pre-JWST data, including the 2017 VLT/Ks measurement at 2.1 m, can be well-fitted with mixtures dominated by micrometre-sized water ice or silicate grains with the contribution of very small grains, also indicating a nearly constant opacity throughout the near-infrared. However, when the JWST data are included, it is not possible to find a solution (i.e., a combination of two materials and grain sizes) that simultaneously fits all datasets, particularly the JWST F150W2, F322W2, and the VLT/Ks bands. The best-fit curves in this case contain predominantly silicates with 1–10 m grain sizes, but the overall fits are poor. This clearly indicates that the observed change in opacity cannot be solely attributed to wavelength dependence, and leaves us with the most likely scenario that the opacities in C1R have indeed increased.
The small particle sizes inferred for C2R raise questions about its long-term stability. In general, radiation pressure could expel micron-sized grains on relatively short timescales, potentially leading to structural distortion or dispersal [25]. Yet, occultation data spanning nearly a decade reveal no significant changes in the ring’s radial position (see Table 1), suggesting that a stabilising or replenishing mechanism must be active. Such mechanisms may mirror those proposed for other dusty rings and could involve interactions with C1R. The confinement of C1R is thought to result from the strong 1:3 spin-orbit resonance with Chariklo’s rotation, in which a ring particle completes one orbit for every three rotations of the central body [33]. While effective for C1R, this resonance drives an outward migration of the ring, ultimately risking its departure from the resonance and subsequent dispersion. A small external shepherd moon has been hypothesised to prevent such drift and maintain C1R’s location [33]. If present, this moonlet might share its orbit with C2R and act as a source of dust, analogous to the small satellites observed in Saturn’s and Uranus’s ring systems. This scenario is further supported by our estimate of C2R’s dynamical spreading timescale: based on JWST measurements, we find a median value of only 0.36 years (1 range: 0.23–0.77 years; see Methods, Section 1.6), implying that without confinement or replenishment, the ring should have dispersed since its discovery in 2013. A co-orbiting shepherd satellite could thus play a dual role—maintaining the ring’s sharp edges while continuously supplying fresh material [17]. However, another recent dynamical study about the effect of solar radiation pressure on the stability of a small-body rings also shows that for highly tilted rings, such as around Chariklo, there are two plausible particle size range that can be stable: m and m, continuing to allow for both possible options [18].
The emerging picture suggests that the two scenarios for C2R—rapid erosion and a dusty, micron-scale particle population—are not mutually exclusive. A ring composed of sub-micron grains may be inherently short-lived, as such particles are particularly susceptible to dynamical processes that promote dispersal. We therefore consider both effects jointly, as they may together account for the observed decline in opacity and signal strength of the C2R ring. If the outer ring is indeed degrading over a timescale of just a few years, this would mark it as a transient structure, potentially representative of a broader class of ephemeral rings among small bodies, as hinted by episodic activity and temporary bound structures around another Centaur, Chiron [10]. If instead the C2R ring is linked to Chariklo’s activity as a Centaur-like object, we would again expect small grains, as cometary dust is predominantly sub-micron in size [34, 35]. Crucially, a single occultation in visible light of a sufficiently bright star could distinguish between these scenarios: if the opacity matches values from 2013–2017, grain-size effects are favoured; but if the occultation yields significantly lower opacities or no detection at all, it would instead support the interpretation of an actively fading ring. Either outcome would raise fundamental questions about the longevity of dusty rings and the fate of metre-sized material subject to slow but persistent erosion over decadal timescales. Ultimately, our JWST observations unveil a new and unexpected level of complexity in minor body ring systems—one that challenges existing paradigms and opens an unprecedented window into the transient architectures sculpting the outer Solar System. These findings provide the first evidence of such episodic changes in these seemingly frequent systems, revealing a dynamical behaviour that had remained entirely hidden until now.
| Parameter | F150W2 | F322W2 | ||||
| C1R | C2R | C1R | ||||
| 1st contact | 2nd contact | 1st contact | 2nd contact | 1st contact | 2nd contact | |
| Mid-time | 13:35:19.74 | 13:40:02.72 | 13:35:13.15 | 13:40:09.15 | 13:35:19.64 | 13:40:02.66 |
| [km] | ||||||
| [km] | ||||||
| [km] | ||||||
| [km] | ||||||
Note: : Transmittance; Mid-time: midpoint time of each occultation contact; : radial distance in the ring plane (km); : radial width in the ring plane (km); : apparent opacity; : normal opacity; : apparent optical depth; : normal optical depth; : equivalent width (km); : equivalent depth (km). C2R was not detected in the F322W2 filter due to a low signal-to-noise ratio. 1st and 2nd contact refer to before and after the closest approach, respectively. Values are presented with uncertainties: asymmetric uncertainties are indicated with upper and lower limits (), while symmetric uncertainties are shown with a single value following the symbol.
1 Methods
1.1 Predicting and observing stellar occultations from space
In June 2020, the first-ever predicted stellar occultation observation using a space telescope was obtained with the Characterising ExOPlanets Satellite (CHEOPS) [36]. This observation demonstrated the feasibility of predicting and observing stellar occultations caused by distant solar system minor bodies using space telescopes: these facilities are unaffected by weather conditions and can extend the range of observations beyond the Earth’s surface. The use of space telescopes also facilitates the detection of material surrounding small bodies, as it enables photometric accuracy comparable to that achieved by ground-based telescopes with much larger apertures. We also have access to stellar occultations of fainter stars than from the ground, multiplying the number of possible events.
JWST is orbiting close to the second Sun-Earth Lagrange point (L2) [37, 38]. L2 is an unstable equilibrium point, meaning constant manoeuvres must be made to keep the telescope in orbit around it [39]: such manoeuvres are usually made 2–3 times a month. Consequently, the predicted orbital path of JWST needs to be frequently updated (J. Giorgini, personal communication, June 22, 2022). For all the above reasons and due to the complex motion of the space telescope around L2, predicting stellar occultations by distant minor bodies visible from JWST is a very challenging task. Extended Data Figures shows the difference between two consecutive releases of the JWST ephemeris, obtained on September 27, 2022. In just under six months from the initial epoch, the difference between the two releases reached approximately 10,000 km, which is larger than the Earth’s radius. In the absence of formal positional uncertainty, the plot illustrates the trustworthiness of a stellar occultation prediction using JWST far from the initial epoch.
To predict stellar occultations that JWST could potentially observe, we used the Stellar Occultation Reduction and Analysis (SORA) software package [40] combined with the Gaia DR3 catalogue [41, 42] and the NIMA ephemeris [43] for the solar system bodies. The strategy we developed to run the predictions is as follows. We selected the 40 largest TNOs/Centaurs (36 TNOs and 4 Centaurs), then ran predictions for these 40 targets every week for Gaia DR3 stars up to a G-magnitude of 20. This frequency of re-running the predictions was necessary due to the station-keeping manoeuvres applied to JWST, which changed the position (ephemeris) of the space telescope. Each week, we reviewed the updated predictions to identify favourable events, refined these predictions with new astrometric data obtained from the ground, and reran the predictions for those objects. Finally, when we were confident in the prediction, we activated the target of opportunity (ToO) to observe the occultation from JWST. Since this was a non-disruptive ToO, it had to be activated at least 14 days in advance of the event.
Using the described strategy, we identified the occultation by the ringed Centaur Chariklo on 18 October 2022 (UT), with a low relative velocity (2.5 km/s) and a star bright enough to be observed by JWST (K = 15.7 mag) as the most promising event. We first identified this occultation on 19 August 2022 (blue line in Extended Data Figures). Then we reran the prediction every week to account for changes in JWST’s orbit caused by station-keeping manoeuvres. Our most recent prediction was made on the day of the occultation, just a few hours before the event (green line in Extended Data Figures). In just about two months, the path of the star behind Chariklo, as seen by JWST, moved 110 km, enough to miss the occultation by the main body. Furthermore, the instant of occultation also shifted by 1.5 minutes.
1.2 Observations
We used the James Webb Space Telescope (JWST) to observe a stellar occultation by Chariklo’s rings, marking the first time such an event was both predicted and successfully observed by JWST. It also represents the first stellar occultation by a small Solar System body ever recorded from beyond Earth’s orbit, with the only comparable case being a previous Quaoar occultation observed by the Earth-orbiting CHEOPS satellite. This observation was both scientifically valuable and technically demanding. It was conducted under JWST’s Solar System Guaranteed Time Observations (GTO) program #1271, ToO TNOs: Unveiling the Kuiper belt by stellar occultations (PI: Pablo Santos-Sanz) [44].
On 18 October 2022, between 12:36:04 and 13:50:27 UT, we observed the event with the JWST NIRCam instrument, targeting the star Gaia DR3 6873519665992128512 (RA 20h 21m 46s, Dec -16∘ 26′ 22′′; G = 17.5, K = 15.7). Observations were performed simultaneously using the wide F150W2 (0.6–2.3 m) and F322W2 (2.4–5.0 m) filters to maximise the stellar flux, a key factor given the high-cadence time series required.
We used the smallest available NIRCam subarray, SUB64P (64 64 pixels; 2′′ 2′′ and 4′′ 4′′ fields of view for the short- and long-wavelength channels, respectively), in combination with the RAPID readout pattern. Each integration comprised five non-destructive reads (groups), with a total exposure time of 0.304 s and a dead time of only 0.0009 s between integrations, resulting in a cadence of 3.3 Hz. At Chariklo’s distance (17 au), this cadence corresponds to a spatial resolution of 750 m per integration. In total, 14,827 integrations were acquired over the 1.25-hour observation window. All data were processed using JWST calibration pipeline version 1.10.0 and the reference file context jwst_1075.pmap.
1.3 Photometry: occultation light curves
Aperture and point spread function (PSF) fitting photometry were performed independently on the JWST/NIRCam time-series data obtained with the F150W2 and F322W2 filters. Aperture photometry yielded a lower dispersion than PSF-fitting methods and was thus adopted for the occultation analysis. Given the minimal field of view (FoV) of the SUB64P subarray, no suitable reference stars were available for differential photometry; only Chariklo and the occulted star were present in the images (see left Panel in Extended Data Figures). The aperture radius was optimised to maximise the signal-to-noise ratio (SNR) of the combined flux from Chariklo and the star, while minimising photometric scatter. Fluxes were normalised to unity outside the occultation events.
The analysis focused on a subset of 1,646 images (out of 14,827 total exposures), centred around the predicted occultation time. The measured flux dispersion was 0.025 in the F150W2 filter and 0.041 in F322W2. A small fraction of the images—approximately 2% in F150W2 and 1.2% in F322W2—were affected by unidentified artefacts and excluded from the analysis.
The resulting light curves (Figure 1) display the drop in stellar flux as Chariklo’s rings transited the background star. As anticipated from the occultation geometry (Extended Data Figures), the main body of Chariklo did not intercept the line of sight, and no additional dips beyond the 2 noise level were detected, even when the complete data set was examined.
Fluxes from Chariklo and the target star were also measured separately in both filters, using images acquired before the occultation when the two sources were spatially resolved (Extended Data Figures, middle Panel). These measurements were used to calibrate the relative contributions of Chariklo and the star to the combined flux, a crucial step in deriving the ring parameters discussed in Section 1.4. The fractional flux contribution of Chariklo (relative to the combined Chariklo + star flux) was 0.2780 for F150W2 and 0.2662 for F322W2.
1.4 Analysis of the rings
In the obtained light curves, the main, densest, and most opaque ring of Chariklo is clearly visible. This ring is called C1R, following the original notation [7]. The C1R occultation light curve displays obvious diffraction spike features immediately before and after the ingress and egress, respectively. These spikes provide strong evidence that the ring has sharp edges and is tightly confined, similar to the rings of Uranus [23, 24]. However, the dynamical mechanism responsible for this confinement is still unknown and is currently under discussion [16, 45, 46, 33].
The other ring (referred to as C2R) is surprisingly not clearly detected in the F150W2 or F322W2 light curves, as shown in Extended Data Figures. Since a ring partially occults the starlight, we fit an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction, finite bandwidth, exposure time, and the Gaia DR3 stellar diameter. We use the procedures within the SORA package [40]. From this analysis, we obtained the time when the star disappeared and reappeared behind Chariklo’s rings ( and ) and the observed apparent ring opacity (), which measures the fractional drop of stellar flux.
Knowing the reconstructed JWST position at the time of the occultation, we were able to propagate the instants of disappearance and reappearance of the star into the tangent plane and determine two-dimensional coordinates (, ) of the rings’ positions relative to Chariklo’s expected centre. This projection was executed using the SORA package, and its result is shown in Figure 1. Assuming Chariklo’s parameters (body and rings) as published by [15], we fitted the centre position of Chariklo as viewed by JWST at the mid-point of the occultation (18 October 2022 at 13:37:42.18 UT), which resulted in
That meant an appulse with the closest approach of 132.2 km, in contrast to Chariklo’s volume-equivalent radius of 124.8 km [15].
Initially, the parameters of both rings (C1R and C2R) were obtained in the sky plane simultaneously, just before and after the closest approach. We used the formalism presented by [14] to project these parameters to the ring plane so we could obtain the relevant parameters, such as the ring radial width (), the normal opacity (), and the normal optical depth ().
Assuming a monolayer ring, the observed ring opacity normal to the ring () can be described as
alternatively, if we assume a poly layer ring, is defined as [47]
where is the ring opening angle, and (resp. ) corresponds to a pole-on (resp. edge-on) viewing. On the other hand, considering that the Airy scale is larger than the width of the ring as seen in the sky plane [48], a factor is needed to compute the fraction of lost light; in that case
for a monolayer or polylayer ring, respectively.
The integrals of and over the radial width () of the ring define the equivalent width () and the equivalent depth () of the profile. These values ( and ) are proportional to the amount of material present in the profiles for each case [14]. The ring parameters are obtained in Table 1. We highlight that the detection of C2R is 1 significant in the F150W2 data and cannot be detected in the F322W2 data.
1.5 Sampling a ring with azimuthally variable structure
In the case of a ring with an azimuthally variable structure, we may observe different normalized opacities () or equivalent with () values at the 1 and 2 contacts during an occultation event, as well as variations between multiple occultations. If the ring opacity is assumed to be wavelength-independent, the observed decrease/increase in opacity and/or in during the 2022 JWST events—relative to previous occultations—could be explained by the possibility that both ingress and egress sampled low/high-density segments of the ring. To estimate the likelihood of such an occurrence, we performed a simple Monte Carlo simulation. We modelled an azimuthally variable ring, with the material distribution given by or , where is the azimuthal angle and is a small integer. Random chords drawn through this ring yield a distribution of ‘observed’ or values.
In the real occultation measurements of the C2R ring, we obtain in seven of the occultations (see Extended Data Tables), with significantly lower values observed only during the JWST event. For simplicity, we assume that the minimum of along the ring is zero, although a constant positive offset may exist. Therefore, the probability of randomly sampling a pair of such low values in just one of eight events, as calculated from the simulation, represents an upper limit, with a probability of obtained in our simulation. In a similar way, we obtain a probability of p 0.004, assuming that we sample the ring close to the density maxima in the JWST measurements, and below these values in the pre-JWST cases (Extended Data Tables). While azimuthal fragmentation is plausible—especially given the existence of fragmented rings in the trans-Neptunian region, e.g. in the case of Quaoar [36, 22]—these very low probabilities suggests that the observed JWST opacity drop (C2R) or increase (C1R) is unlikely to be solely due to this effect. The results have been found to be independent of the parameter for 3. Here, we assumed that the azimuthal structures of the two rings are independent.
1.6 Spreading timescale of Chariklo’s C2R Ring
We compute the characteristic spreading timescale of Chariklo’s second ring (C2R) using the standard expression for a narrow, self-gravitating ring in resonance [49]:
where is the gravitational constant, the mass of Chariklo, the ring semi–major axis, and its radial width. Chariklo’s mass is derived from its bulk density and ellipsoidal volume:
We adopted ellipsoidal semi-axes from Morgado et al. (2021)[15]: km, km and km and a mean bulk density of .
From our JWST occultation measurements of the C2R ring’s and contact radii ( and km), we obtain the weighted mean:
To account for the uncertainty in radial width, we uniformly sampled over the KDE-inferred interval [0.40, 3.25] km. A Monte Carlo simulation with realisations yielded a median spreading timescale of s (0.36 yr), with a 1 confidence interval between s (0.23 yr) and s (0.77 yr).
1.7 Radiative Transfer Modeling
After the striking marginal (at F150W2) and non-detection (at F322W2) of the C2R in the JWST filters, wavelength-dependent observables induced by micron- and submicron-sized grains became a real concern in explaining the measured discrepancies against previous occultations. Here, we perform a thorough analysis of the compatibility of dusty ring models for both C1R and C2R. Our main goals are to investigate which material in our models can replicate these observations and what we can say about the possible presence or absence of water ice in the rings. To explore whether a potential wavelength-dependent effect is responsible for the observed changes in the measured opacities, we conducted a radiative transfer analysis, assuming that the ring did not change during the observation period (2013-2022) and that all observational chords sampled the same amount of material, after correcting for the opening angle change due to Chariklo’s orbital motion around the Sun. We used a slab-like, two-component model, where we mixed materials to match their common optical properties with the measured occultation normal opacities () or, equivalently, their normal optical depths (). In our model, the optical depth is obtained as the sum of the optical depths of two materials (two different grain sizes and/or compositions):
which is calculated from the mass densities and , and using the extinction coefficient = , the sum of the absorption and scattering coefficients:
where is the slab’s length (the ring’s thickness). A single-component model is obtained if is chosen. To calculate the extinction coefficients, we used the OPTOOL package [50], which provides direct absorption and scattering coefficients ( and in units of cm2/g) for different types of materials and grain sizes. Previous studies based on reflectance spectra between 1998 and 2013 proposed an icy ring system around Chariklo and found that it may contain 20% water ice and 80% other materials, such as silicates, organics and small quantities of carbon [19], which motivated our choice of materials to investigate: water ice, carbon and silicates. OPTOOL offers both amorphous and crystalline options for most materials, and we found no significant differences between these options for our measurements. Therefore, we used the amorphous option in our analysis. We tested different kinds of silicates, such as pyroxene with various magnesium contents, but found minimal impact on the solutions. Therefore, we used a pyroxene model with 100% magnesium content, referred to as silicates hereafter.
In general, although is a more observationally constrained value, compositional information is carried by the opacity estimates. Model calculations (see Equation 9) can provide only optical depth values for direct comparison. However, in , this information is lost when it is combined with the radial width of the measurements. To compare the observed occultation opacities with our models, we consider the spectral energy distribution of the occulted star and each filter’s transmission curves . For a specific filter , the ratio of occulted and non-occulted photon counts is given by:
where is the wavelength-specific opacity, obtained from the model optical depths as = .
First, we used single-grain models with sizes between 0.2 and 10 m. Sizes larger than 10 m do not show significant differences in extinction properties at different wavelengths in the visible to near-infrared range. We note that for planetary rings, we expect the grain size range to show a power-law distribution with the exponent around three [51]. However, with the accuracy of the occultation measurements available, our primary goal was to identify the most dominant grain sizes. In each case, we determined the best-fit models using least squares minimisation. The values were calculated as:
where = with being the normal measured opacity, and is its uncertainty in the different filters. After identifying the best-fit grain sizes, we mixed the single-grain solutions using two approaches: (1) mixing the same material with two-grain sizes close to each other and (2) mixing two materials with their preferred grain sizes. The cases with the lowest values are considered as our compositional solutions for the rings.
Even though there have been a large number of occultation measurements since 2013, only a few sites could capture the very narrow C2R ring. With their remarkable sensitivity and sampling abilities, the JWST measurements provided a unique opportunity to enrich the existing observations of the secondary ring C2R. Extended Data Tables shows all the available measurements from 2013 on the C2R ring, where the JWST ones are the most well-sampled of all those measurements. Some of the earlier C2R measurements contain only one or two data points, resulting in an unresolved detection of the ring. Therefore, it is essential to have a good understanding of the , and radial width values from previous observations and their associated uncertainties in the probability space.
The most important of all is the VLT measurement in 2017 using the band, which samples the same, albeit smaller, piece of the JWST F150W2 filter range. Panel (a) in Extended Data Figures presents a comparative analysis of Chariklo’s ring system as observed with the VLT in 2017 (at 2.1 m, very close to the JWST F150W2 band) and with JWST in 2022. The radial brightness profiles reveal consistent locations and morphologies of the C1R and C2R ring components across both epochs and instruments, with only minor differences attributable to changes in viewing geometry and filter transmission curves. The JWST observations in the F150W2 and F322W2 filters provide higher signal-to-noise ratios, allowing finer sampling of the ring structure and sharper detection of the main features compared to ground-based data. Notably, however, while a stronger detection of C2R was expected with JWST, the opposite trend is observed: the C2R signal appears significantly weaker, especially in the F322W2 filter, where the z-score of the 1 and 2 contacts reaches only ~0.31 and ~0.23, respectively. Even in F150W2, where the JWST detection is somewhat clearer (~1.70 and ~1.73), it remains below the VLT levels (~4.42 and ~3.36). Nevertheless, the previously published opacity values for this observation appear asymmetrical at first glance. The 1 contact has a low opacity and well-constrained error margin (), which aligns well with the JWST measurements. However, the second contact appears completely different, with considerable uncertainty () -see Extended Data Tables. Figure Extended Data Figures Panel (b) shows the actual distributions of the possible apparent opacity values of this VLT occultation. Below the red lines, the 1- and 3- regions can be found. This is a typical result of fitting the occultation light curves using a Monte Carlo approach with different possible models and evaluating them. For the 1 contact, we can confirm that the value is fairly constrained. Still, for the second one, even though there is (numerically) a best-fitted value (), there is little to no difference between the values within a wide range (which resulted in error bars). This clearly demonstrates that, due to the low resolution of the VLT measurement, we cannot determine a single value for , but rather a range that, as with the 1 contact, overlaps with the JWST F150W2 region.
Because the previously published individual opacity and width values may not accurately represent the most likely values within their given error bars, we used a probabilistic approach to examine the typical radial width values of the C2R ring. To identify the most representative radial widths () derived from the occultation light curve fits, we performed a kernel density estimation (KDE) using all valid C2R measurements. The resulting probability density function, shown in Figure Extended Data Figures, exhibits two distinct peaks that characterise the most probable width regimes ().
These KDE-derived peaks guided the recalculation of the normal optical depth , using the relation:
where is the equivalent width obtained from the light curve fitting.
When the measured of an observation overlapped with a KDE peak, that value was directly used in Equation 12. In contrast, if the measured did not intersect any peak, a constrained search within its asymmetric uncertainty bounds was performed to select the value closest to the nearest KDE maximum, while ensuring it remained within the original asymmetric limits. From the set of candidate values yielding within the observational error limits, the extremal value (either minimum or maximum) was selected based on the direction of the nearest peak. This procedure was applied, for instance, to observations such as OPD (both contacts), Springbok (second contact), and Gifberg (second contact), which do not directly overlap with either KDE peak. This approach avoids arbitrary resampling and preserves both physical consistency and fidelity to observational uncertainties.
By applying this method across all measurements, we derived new, more probable normal optical depth values, , constrained strictly within the original asymmetric uncertainties. The resulting distribution, shown in Figure 3, illustrates the original uncertainty ranges along with the resampled values (Extended Data Tables). The KDE-based adjustment significantly reduces the spread of the data points, highlighting consistent trends that were previously obscured. For instance, the two VLT observations, which originally showed substantial differences in (see Extended Data Tables), now exhibit closer agreement once KDE-informed widths are applied. However, the second, much larger VLT measurement also covers a wider potential opacity range with an overlap with the first, well-constrained value, which also agrees well with the JWST F150W2 observation. Our analyses indicate a reasonable likelihood that the two 2017 VLT measurements exhibit similar opacities to the 2022 JWST measurements, suggesting a silicate-like composition consisting of small grains. Considering also the Danish observations, the preferred silicate solutions for the C2R ring are between 0.2 – 0.5 m. The two Danish and the two JWST observations are crucial because, for each instrument, the observations in two filters were parallel in time and sampled the same part of the ring. Therefore, they are not sensitive to possible inhomogeneities in the ring.
The bottom Panel of Figure 3 presents the comparison between the resampled (and original) observations and the models. The normal opacities transformed into relative intensities as , showing the fractional dimming of the occulted star, which now can be directly compared to the models. For the non-detection in the F322W2 filter, we indicated a violet range set to a relative intensity of 1.0, where the width of the region represents the uncertainty of the non-detection – larger signals should have been detected. We created pyroxene models with a magnesium content of 100% using 0.2, 0.3, 0.5, 1.0 and 3.0 m grain sizes, marked as red lines in the bottom panel of Figure 3. These silicate models can capture the general trend of the C2R opacities, which suggests that opacities in the visible wavelength are more prominent than the near-infrared ones. This decreasing signal strength aligns well with the fact that the smaller the grain size of the silicate models, the greater their observable opacity in the visible range. Therefore, we suggest a compatible dusty ring model for C2R dominated by silicate grains in the range of m. We also note that the optical properties of the pyroxene models are also suitable for explaining an asymmetric detection in the two JWST filters, as based on these models, we expect a higher signal in the F150W2 compared to F322W2. Since the JWST measurements in these filters were taken at the same time, sampling the same ring segment, such a silicate-dominated dusty ring may also contribute to an unexpected difference between the two filters. All these models show that in the F322W2 band, no significant signal can be detected, if grain sizes are kept below 3 m, as represented by the red curves confined to the violet-coloured region.
The sky blue lines correspond to the water models with grain sizes of 0.5, 1.0 and 4.0 m, which display a distinctive signature feature at around 3 m. This particular feature is revisited in Figure 4, where we mainly focus on the JWST filter ranges. The middle panel of Figure 4 clearly shows that the smaller the grain size, the larger the feature itself. For grains larger than 2-5 in size, the curves flatten, distributing nearly equal contribution among all wavelengths. Fortunately, the wavelength range of this water signature is entirely covered by the F322W2 passband, offering a natural selection between the suggested models in this scenario. Taking the transmission profiles of each filter into account (top Panel in Figure 4), water ice grains of m would make a notable contribution in the F322W2 filter, relative to the F150W2 one. Pure water ice rings, similar to those around Saturn [30], cannot explain the measurements for either sub-micron or larger particles. Although our numerically best fit model for all the data is clearly the pure pyroxene one (see Extended Data Figures), the finite resolution of the measurements leaves room for other, similarly suitable mixture models. Therefore, the presence of a certain amount of water ice remains compatible with the measurements. However, in this purely wavelength-dependent scenario, the exact amount of water ice in C2R heavily depends on the chosen best-fit silicate model and the assumed water grain sizes. The bottom panel of Figure 4 exhibits silicate models with grain sizes of m, combined with different amounts of m, (solid) or m (dotted) ice grains. Besides the exact differences, what is the same in all models is that water content above 10-30% would lead to a decrease in the observed visual-infrared opacity discrepancy and a significant increase in the contribution of the F322W2 filter, which would then have been observed along with the marginal detection in F150W2. These estimates of water content are also very similar to those previously suggested for the C1R ring [19].
To investigate the wavelength-dependent opacity of C1R, we performed a similar, simple radiative transfer calculation assuming a homogeneous layer for each occultation event, composed of a mixture of two materials with different compositions and grain sizes. The pre-JWST data, including the 2017 VLT/Ks measurement at , can be well fitted with a mixture of silicates (olivine or pyroxene) with grain sizes and water ice with grain sizes between 1–10 (blue curve in Figure 5). This fit implies nearly constant opacity across the near-infrared, with increased opacity at shorter visible wavelengths. Comparable fits can also be achieved using mixtures of silicate grains with two distinct size populations, similar to one above. Thus, these data do not strongly constrain the material composition of C1R. However, grains larger than would produce nearly wavelength-independent opacity over the full range considered, regardless of composition, suggesting that the grains responsible for most of the extinction in C1R are likely smaller than this threshold. When the JWST data are included, however, no combination of two materials and grain sizes provides a satisfactory fit across all datasets—particularly the JWST F150W2, F322W2, and VLT/Ks bands. The best-fit models in this case (red curve in Figure 5) require predominantly silicates with 1–10 grain sizes, but the overall fit quality remains poor. This strongly suggests that the observed opacity variations cannot be explained solely by wavelength dependence. For comparison, spectroscopy of the unresolved Chariklo system indicates a composition of 20% water ice mixed with darker materials, probably silicates [19].
Extended Data Figures
Extended Data Tables
| Telescope/Instrument/Filter | Date |
| Danish-1.54m/DFOSC/Za | 2013-06-03 | ||||||
| VLT/Hawk-I/Ksa | 2014-02-16 | ||||||
| SAAO/SHOCa | 2014-04-29 | ||||||
| Springbok/Raptor/Emptya | 2014-04-29 | ||||||
| Hakos/Raptor/Clearb | 2017-04-09 | - | |||||
| WabiLodge/Raptor/Clearb | 2017-04-09 | ||||||
| WeaverRL/Raptor/Clearb | 2017-04-09 | ||||||
| Tivoli/Raptor/Clearb | 2017-06-22 | ||||||
| WdhBackes/Raptor/Clearb | 2017-06-22 | ||||||
| Onduruquea/Raptor/Clearb | 2017-06-22 | ||||||
| Outeniqua/Raptor/Clearb | 2017-06-22 | ||||||
| WdhMeza/ASI178MM/Clearb | 2017-06-22 | ||||||
| Hakos/Raptor/Clearb | 2017-06-22 | ||||||
| Danish-1.54m/Redb | 2017-07-23 | ||||||
| VLT/Hawk-I/Ksb | 2017-07-23 | ||||||
| SSO/Andor/g/I+zb | 2017-07-23 | ||||||
| ElRodeo/MerlinKite/Clearb | 2017-07-23 | ||||||
| ElSalvador/Raptor/Clearb | 2017-07-23 | ||||||
| Danish-1.54m/Visualb | 2017-07-23 | ||||||
| LaSilla1m/Raptor/Clearb | 2017-07-23 | ||||||
| OPD/Andor/Clearb | 2017-07-23 | ||||||
| TolarGrande/Raptor/Clearb | 2017-07-23 | ||||||
| IncaDeOro/Raptor/Clearb | 2017-07-23 | ||||||
| OPD/Andor/Clearb | 2017-08-24 | ||||||
| Glenlee/Watec/Clearb | 2020-06-19 | ||||||
| JWST/NIRCam/F150W2c | 2022-10-18 | ||||||
| JWST/NIRCam/F322W2c | 2022-10-18 |
| Telescope/Instrument/Filter | Date | Ring | |||
| Danish-1.54m/DFOSC/Za | 03/06/2013 | C1R | |||
| Springbok/Raptor/Emptya | 29/04/2014 | C1R | |||
| Danish-1.54m/Redb | 23/07/2017 | C1R | |||
| Danish-1.54m/Visualb | 23/07/2017 | C1R | |||
| OPD/Andor/Clearb | 23/07/2017 | C1R | |||
| VLT/Hawk-I/Kb | 23/07/2017 | C1R | |||
| JWST/NIRCam/F150W2c | 18/10/2022 | C1R | |||
| JWST/NIRCam/F322W2c | 18/10/2022 | C1R |
| Telescope/Instrument/Filter | Date | |||||||
| (m) | (km) | (km) | (km) | (km) | ||||
| Danish-1.54m/DFOSC/Za | 03/06/2013 | 1.054 | ||||||
| Springbok/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Gifberg/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Danish-1.54m/Visualb | 23/07/2017 | 0.510 | ||||||
| Danish-1.54m/Redb | 23/07/2017 | 0.770 | ||||||
| OPD/Andor/Clearb | 23/07/2017 | 0.640 | ||||||
| VLT/Hawk-I/K | 23/07/2017 | 2.145 | ||||||
| JWST/NIRCam/F150W2c | 18/10/2022 | 1.655 | ||||||
| JWST/NIRCam/F322W2c | 18/10/2022 | 3.225 | - | - | - | - |
| Telescope/Instrument/Filter | Date | |||||||
| (m) | (km) | (km) | (km) | (km) | ||||
| Danish-1.54m/DFOSC/Za | 03/06/2013 | 1.054 | ||||||
| Springbok/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Gifberg/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
| Danish-1.54m/Visualb | 23/07/2017 | 0.510 | ||||||
| Danish-1.54m/Redb | 23/07/2017 | 0.770 | ||||||
| OPD/Andor/Clearb | 23/07/2017 | 0.640 | ||||||
| VLT/Hawk-I/Ksb | 23/07/2017 | 2.145 | ||||||
| JWST/NIRCam/F150W2c | 18/10/2022 | 1.655 | ||||||
| JWST/NIRCam/F322W2c | 18/10/2022 | 3.225 |
Acknowledgements
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1271. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. P.S-S. acknowledges financial support from the Spanish I+D+i project PID2022-139555NB-I00 (TNO-JWST) funded by MCIN/AEI/10.13039/501100011033. Y.K. is forever thankful to TÜBİTAK National Observatory (1997-2024) for its unwavering support of his career and science. The authors acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MICIU/AEI/ 10.13039/501100011033. M.A. thanks CNPq grants 427700/2018-3, 310683/2017-3 and 473002/2013-2. A.R.G.J. thanks the financial support of São Paulo Research Foundation (FAPESP, proc. 2016/24561-0 and 2018/11239-8). C.L.P. is thankful for the support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro - FAPERJ/DSC-10 (E26/204.141/2022). HBH and SNM also acknowledge support from NASA JWST Interdisciplinary Scientist grant 21-SMDSS21-0013.

### v2

@@
 Abstract
-Ring systems have recently been discovered around several small bodies in the outer Solar System through stellar occultations.
-While such measurements provide key information on ring geometry and dynamical interactions, little is known about their composition, grain-size distribution, origin, lifetime, or evolutionary pathways.
-Here we report near-infrared observations from the James Webb Space Telescope (JWST) of a stellar occultation by the Centaur (10199) Chariklo, providing unprecedented constraints on the material properties of a small-body ring system and offering insights into their origin and evolution.
-These measurements reveal that Chariklo’s inner dense ring contains predominantly micrometre-sized particles and exhibits a significant increase in opacity compared to previous observations, suggesting active replenishment events.
-Most strikingly, the outer ring shows a much weaker near-infrared occultation signature than in earlier visible-light detections.
-This discrepancy may indicate ongoing material loss, implying that the outer ring is transient, or it may reflect wavelength-dependent opacity consistent with a dusty structure dominated by 0.2–0.5 m silicate grains.
-These scenarios, not mutually exclusive, point to an unprecedented level of complexity in small-body ring systems, unlike anything observed around other minor bodies in the Solar System.
+Ring systems have been discovered around several small bodies in the outer Solar System through stellar occultations.
+While such measurements provide key information about ring geometry and dynamical interactions, little is known about their origins, lifetimes, evolutionary pathways, or compositions.
+Here we report near-infrared observations with the James Webb Space Telescope (JWST) of a stellar occultation by (10199) Chariklo, a Centaur known to host a double-ring system.
+Our JWST measurements show that Chariklo’s inner dense ring has become significantly more opaque than in previous observations, pointing to ongoing replenishment processes or dynamical restructuring.
+In contrast, the outer ring exhibits a much weaker near-infrared occultation signature than seen in earlier visible-light detections.
+This discrepancy may reflect material loss, suggesting that the outer ring could be transient, or may arise from wavelength-dependent opacity.
+These scenarios, which are not mutually exclusive, point to an unprecedented level of complexity in small-body ring systems, distinct from those observed around any other minor bodies in the Solar System.
 keywords:
 KBOs, TNOs, Centaurs, stellar occultations, JWST, Chariklo, ringsStellar occultations are a powerful tool in planetary astronomy, offering unique opportunities to probe the properties of distant solar system bodies and their environments with exceptional precision.
-When a celestial object, such as an asteroid, passes in front of a star, the brief dimming of starlight reveals information about the occulting body’s size, shape, and surroundings [1, 2].
-This technique has also enabled the detection and characterisation of tenuous atmospheres [3, 4, 5], or has placed stringent upper limits on their presence [2].
-Notably, stellar occultations have led to the discovery and study of ring systems not only around the giant planets [6], but more recently around small bodies, including the Centaurs Chariklo [7] and Chiron [8, 9, 10], the dwarf planet Haumea [11], and the trans-Neptunian object Quaoar [12, 13].
-Among these, Chariklo was the first small Solar System body discovered to host a ring system, composed of two sharply confined components, C1R and C2R, orbiting at 390 and 405 km from its centre, respectively [7, 14, 15].
-Yet, despite extensive occultation campaigns and multiple studies on ring dynamics [16, 17, 18], the composition, grain-size distribution and origin of the rings have remained poorly constrained and have defied a clear explanation for over a decade.
-Constraining the composition of rings around small Solar System bodies remains challenging, as their proximity and faintness preclude the isolation of ring spectra using current remote spectroscopy [19].
-However, wavelength-dependent optical properties of the grains, which lead to different occultation opacities in the visible and near-infrared, may provide direct constraints on grain properties [20], as it has been proposed for Haumea’s ring system [21].
+When a celestial object, such as an asteroid, passes in front of a star, the brief dimming of starlight reveals information about the occulting body’s size, shape, and surroundings [Sicardy2011, Sicardy2006].
+This technique has also enabled the detection and characterisation of tenuous atmospheres [Sicardy2016, Sicardy2021, MarquesOliveira2022], or has placed stringent upper limits on their presence [Sicardy2006].
+Notably, stellar occultations have led to the discovery and study of ring systems not only around the giant planets [Sicardy2024], but more recently around small bodies, including the Centaurs Chariklo [Braga-Ribas2014] and Chiron [Ortiz2015, Sickafoose2020, Ortiz2023], the dwarf planet Haumea [Ortiz2017], and the trans-Neptunian object Quaoar [Morgado2023, Pereira2023].
+Among these, Chariklo was the first small Solar System body discovered to host a ring system, composed of two sharply confined components, C1R and C2R, orbiting at 390 and 405 km from its centre, respectively [Braga-Ribas2014, Berard2017, Morgado2021].
+Yet, despite extensive occultation campaigns and multiple studies on ring dynamics [Sicardy2019, Sickafoose2024, Regaly2025], the origin, composition and grain-size distribution of the rings have remained poorly constrained and have defied a clear explanation for over a decade.
+Although wavelength-dependent optical properties of the grains, which lead to different occultation opacities in the visible and near-infrared, may provide direct constraints on grain properties [bohren2008absorption, Kalup2024], due to the limited number of well-resolved observations and the possible structural changes in the ring, it still remains challenging.
 To investigate the fine structure and composition of Chariklo’s rings beyond the limitations of ground-based observations, we targeted a stellar occultation using the James Webb Space Telescope (JWST).
@@
 This observation was conducted simultaneously in two near-infrared bands, centred at 1.5 and 3.2 m (F150W2 and F322W2), including the first occultation detection ever beyond 3 m, which is impossible from the ground.
-Exactly the same JWST setup was used later in the observations of the occultation event by the ring system of Quaoar [22].
+Exactly the same JWST setup was used later in the observations of the occultation event by the ring system of Quaoar [Proudfoot2025].
 Although the radial locations of both Chariklo’s rings in 2022 remain consistent with earlier measurements, they also exhibit significant variations compared to previous observations.
-For clarity, we adopt the terms first contact ( contact) and second contact ( contact) to refer to the two ring-plane crossings, each comprising an ingress and an egress: Ingress1–Egress1 and Ingress2–Egress2, respectively.
-This terminology will be used throughout the manuscript.
-Chariklo’s inner dense ring, C1R, is clearly detected in the JWST light curves, exhibiting pronounced diffraction spikes at ingress and egress (see Figure 1), indicative of sharp edges and strong confinement, morphological features reminiscent of Uranus’s rings [23, 24].
-Interestingly, C1R is found to be more opaque than before: the pre-JWST measurements provide an average normalized opacity of for this ring, while considering only the JWST measurements gives , indicating a highly significant difference, corresponding to a z-score of z = 4.2 and probability of .
-In contrast, the outer ring, C2R, displays an unexpectedly weak signal in the JWST measurements.
+For clarity, we adopt the terms first contact ( contact) and second contact ( contact) to refer to the two points where the star’s line of sight intersects the ring plane, each comprising an ingress and an egress: Ingress1–Egress1 and Ingress2–Egress2, respectively.
+This terminology will be used throughout the paper.
+Chariklo’s inner dense ring, C1R, is clearly detected in the JWST light curves, exhibiting pronounced diffraction spikes at ingress and egress (see Figure 1), indicative of sharp edges and strong confinement, morphological features reminiscent of Uranus’s rings [Elliot1977, Elliot1987].
+Interestingly, C1R is found to be more opaque than before—with all uncertainties throughout this paper reported at the 1- level—the pre-JWST measurements provide an average normal opacity of for this ring, while considering only the JWST measurements gives , indicating a highly significant difference, corresponding to a z-score of z = 4.2 and probability of .
+In contrast, the outer ring (C2R) shows an unexpectedly weak signal in JWST measurements.
 After applying an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction (see Section 1.4), C2R was marginally detected at 1 in the F150W2 data, but not in the F322W2 filter, despite both measurements being performed simultaneously and sampling the same ring segment.
-Compared to previous ground-based visible range observations, this suggests particularly low opacities for C2R.
+Compared with previous ground-based visible-range observations, this suggests particularly low opacities for C2R.
 To unravel the origin of these differing opacities observed across epochs and wavelengths, we explore three basic scenarios that could explain these phenomena: 1) Azimuthally variable rings, where JWST consistently sampled a sparser or denser segment compared to earlier measurements; 2) Changes in the ring material, involving an increase or decrease in the total amount of material and/or the effective cross-section of grains, but without wavelength dependent grain properties; 3) Grains with wavelength-dependent optical properties, where composition and grain size effects are responsible for the observed variations.
-In the case of a fragmented ring, alternating low- and high-density segments produce azimuthal variability in the observed ring structure, affecting both opacity and equivalent width.
-While previous occultation measurements yield roughly consistent values for these parameters, the JWST observations reveal significantly different opacity and equivalent width values.
-Using a simple azimuthal ring density model (see Section 1.5), we estimate that the probability of sampling only low-density (C2R) or high-density (C1R) segments during both ingress and egress in the JWST observations is very low, with p0.004 for C1R and p0.002 for C2R.
-Changes in the effective cross-section in occultation measurement can be best described by the equivalent width (), which combines opacity, radial width, and the sensitivity and time resolution of the instrument.
-is proportional to the amount of material responsible for the extinction within the occulted segment (see Section 1.4 for its formal definition).
-We carried out a dedicated study of the C2R ring, combining the 2022 JWST occultation measurements (e.g., opacities, equivalent widths) with all available ground-based data from 2013, 2014, and 2017 (see Extended Data Tables and Section 1.4 for details).
-While no clear trend is seen in between 2013 and 2017, a marked decrease is observed between 2017 and 2022 for C2R (Figure 2, Panel (b)), potentially revealing a decade-long temporal evolution.
-This apparent fading of the ring may indicate that C2R is actively dispersing.
+In the case of a fragmented ring, alternating low- and high-density segments produce azimuthal variability in the observed ring structure.
+This affects both opacity and equivalent width (), the latter of which combines opacity, radial width, and the instrument’s sensitivity and time resolution (see Section 1.4 for more information).
+While previous occultation measurements yield roughly consistent values for these parameters, the JWST observations reveal significantly different opacities and equivalent widths.
+Using a simple azimuthal ring density model, we estimate that the probability of randomly sampling a pair of (both ingress and egress) high-density (C1R) or low-density (C2R) values in the JWST observations against eight other events is very low, with p0.004 for C1R and p0.002 for C2R.
+See a comprehensive description of this model and calculation in Section 1.5.
+Changes in the effective cross-section in occultation measurement are best described by the just-cited equivalent width ().
+is proportional to the integrated effective cross-section of the grains, responsible for the extinction within the occulted segment (see Section 1.4 for its formal definition).
+Combining the 2022 JWST occultation measurements with available ground-based data from 2013, 2014, and 2017 (see Extended Data Tables, Extended Data Tables and Section 1.4 for details), we carried out a detailed study of the temporal evolution of the rings.
+In the case of C2R, while no clear trend is seen in between 2013 and 2017, a marked decrease () is observed between 2017 and 2022 (Figure 2, Panel (b)), potentially revealing a fast evolution in a timescale of a few years.
+Strikingly, the JWST opacity and equivalent width of C1R shows an opposite trend, a considerable increase (50%, see Panel (a) in Figure 2, Extended Data Tables, and also [Berard2017, Morgado2021]).
+This change in in C1R () is approximately 10 times greater than that in C2R (), indicating that the material missing from C2R cannot account for the increase in the normal opacity and in C1R.
+This increase could either be due to an overall increase in the amount of material or to the rise in the effective cross-section of grains, for example, through collisions between larger particles or rocks inside the ring, producing a cascade of smaller grains.
+The apparent fading of C2R may indicate that it is actively dispersing.
 Such depletion could result from collisions among particles, producing smaller grains that are more easily lost.
-Without effective confinement, these grains are subject to increasing orbital eccentricities due to radiation pressure, leading their orbits to intersect the inner ring [25].
+Without effective confinement, these grains are subject to increasing orbital eccentricities due to radiation pressure, leading their orbits to intersect the inner ring [Csaba2024].
 Several known dusty rings in the Solar System exhibit temporal variability in opacity, morphology, or confinement, such as Saturn’s D, E, and F rings, Neptune’s Adams arcs, and Uranus’s ring.
-For instance, Saturn’s D ring has shown measurable radial shrinkage and brightness changes on decadal timescales, likely driven by micrometeoroid bombardment or electromagnetic forces [26].
-The E ring, sourced by Enceladus’s plumes, varies with the moon’s activity and solar cycle [27].
-Neptune’s Adams arcs display pronounced morphological evolution and changing brightness over months to years, pointing to a dynamically unstable configuration [28].
-Uranus’s ring also shows opacity changes possibly linked to variable confinement mechanisms [29].
+For instance, Saturn’s D ring has shown measurable radial shrinkage and brightness changes on decadal timescales, likely driven by micrometeoroid bombardment or electromagnetic forces [HEDMAN200789].
+The E ring, sourced by Enceladus’s plumes, varies with the moon’s activity and the surrounding magnetospheric environment.
+[KEMPF2008420].
+Neptune’s Adams arcs exhibit pronounced morphological evolution and brightness changes over months to years, suggesting a dynamically evolving configuration [renner2014].
+Uranus’s ring also shows opacity changes possibly linked to variable confinement mechanisms [FRENCH2012181].
 These examples illustrate that dusty, low-mass rings can undergo rapid and significant evolution, reinforcing the interpretation of Chariklo’s outer ring as a transient, evolving structure.
-Strikingly, the JWST opacity and equivalent width of C1R shows an opposite trend, a considerable increase (see Panel (a) in Figure 2, Extended Data Tables, and also [14, 15]).
-However, the change of in C1R () is approximately 10 times greater than that in C2R (), leaving further questions on the origin of the increase.
-In C1R, this could be due either to an overall increase in the amount of material or to the rise in the effective cross-section of grains, for example, through collisions between larger particles or rocks inside the ring, producing a cascade of smaller grains.
 Since JWST probes a wavelength range where stellar occultations have rarely been observed, the discrepancy with visible-light data may also reflect the optical properties of the ring particles—specifically, a dominance of sub-micron grains whose scattering efficiency decreases at longer wavelengths.
-Even in the pre-JWST observations, a possible sign of wavelength dependence is the significantly different opacities of the 2013 and 2017 occultation measurements of C2R, using the Danish Z (1 m) and Danish Visual (0.55 m) bands (see Figure 3).
-To assess whether these variations reflect a genuine wavelength dependence, we adopt a radiative transfer approach using various grain types (e.g., water ice, olivine, carbon) and particle sizes, and compare the resulting extinction values with all available C2R occultation data (see Extended Data Tables and Extended Data Tables) to constrain plausible compositions consistent with the observations.
+To assess whether these variations reflect a genuine wavelength dependence, we adopt a radiative transfer approach using various grain types (e.g., water ice, olivine, carbon) and particle sizes, and compare the resulting extinction values with available well-resolved occultation measurements (see Extended Data Tables) to provide a compatibility analysis of any plausible compositions that may be consistent with the observations.
 A detailed description of this modelling can be found in Section 1.7.
-We find that for C2R, a wavelength dependence driven by different optical properties of the ring particles cannot be ruled out.
-Models dominated by silicates over water reproduce the detection in the F150W2 filter and the non-detection in the F322W2 filter, with pyroxene-rich compositions fitting both the visible and near-infrared opacities when grain sizes lie between 0.2 and 0.5 m (Figure 3).
-However, we emphasise that these compositional inferences depend on the assumed grain-size distribution and should not be viewed as uniquely limiting the ring’s material composition.
-In any case, if C2R is indeed dominated by micron-sized dust, its population may be analogous to that of Uranus’s ring, which consists predominantly of similar grains [30].
-The ring likely originates from collisions involving a moonlet or a belt of larger parent bodies, potentially metres to kilometres in size, that generate fine particulate debris [31].
-Similar dusty ring structures are seen co-orbiting with small Saturnian moons such as Janus, Anthe, Methone, and Pallene, where dust is believed to originate either from the satellites themselves or from co-orbiting larger bodies [32].
-A comparable dusty ring structure has also been proposed around Makemake [25].
-To explore the wavelength-dependent opacity scenario for C1R, we used a similar radiative transfer calculation, composed of a mixture of two materials with different compositions and potentially different grain sizes (see Section 1.7 and Figure 5).
-The pre-JWST data, including the 2017 VLT/Ks measurement at 2.1 m, can be well-fitted with mixtures dominated by micrometre-sized water ice or silicate grains with the contribution of very small grains, also indicating a nearly constant opacity throughout the near-infrared.
-However, when the JWST data are included, it is not possible to find a solution (i.e., a combination of two materials and grain sizes) that simultaneously fits all datasets, particularly the JWST F150W2, F322W2, and the VLT/Ks bands.
-The best-fit curves in this case contain predominantly silicates with 1–10 m grain sizes, but the overall fits are poor.
-This clearly indicates that the observed change in opacity cannot be solely attributed to wavelength dependence, and leaves us with the most likely scenario that the opacities in C1R have indeed increased.
-The small particle sizes inferred for C2R raise questions about its long-term stability.
-In general, radiation pressure could expel micron-sized grains on relatively short timescales, potentially leading to structural distortion or dispersal [25].
-Yet, occultation data spanning nearly a decade reveal no significant changes in the ring’s radial position (see Table 1), suggesting that a stabilising or replenishing mechanism must be active.
-Such mechanisms may mirror those proposed for other dusty rings and could involve interactions with C1R.
-The confinement of C1R is thought to result from the strong 1:3 spin-orbit resonance with Chariklo’s rotation, in which a ring particle completes one orbit for every three rotations of the central body [33].
+For C1R, we used a mixture of two materials with different compositions and potentially different grain sizes (see Section 1.7 and Figure 3).
+The pre-JWST data, including the 2017 VLT/Ks measurement at 2.1 m, can be well-fitted with mixtures dominated by micrometre-sized water ice or silicate grains with the contribution of tiny grains, also indicating a nearly constant opacity throughout the near-infrared.
+However, when JWST data are included, no solution (i.e., a combination of two materials and grain sizes) can simultaneously fit all datasets, particularly the JWST F150W2, F322W2, and VLT/Ks bands.
+The best-fit curves in this case consist predominantly of silicates with grain sizes of 1–10 m, but the overall fits are poor.
+This clearly indicates that the observed change in opacity cannot be solely attributed to wavelength dependence and leaves us with the most likely scenario: the opacities in C1R have indeed increased.
+For C2R, it is much harder to perform such an analysis due to the lack of well-resolved occultation profiles.
+However, as we introduced above, if C2R is actively depleting, this is a strong indication that small grains dominate this ring, as sub-micron grains may be inherently short-lived and particularly susceptible to dynamical processes that promote dispersal.
+However, small particle sizes raise questions about long-term stability.
+In general, radiation pressure could expel micron-sized grains on relatively short timescales, potentially leading to structural distortion or dispersal [Csaba2024].
+Yet, occultation data spanning nearly a decade reveal no significant changes in the radial position of both C1R and C2R (see Table 1), suggesting that a stabilising or replenishing mechanism must be active.
+The confinement of C1R is thought to result from the strong 1:3 spin-orbit resonance with Chariklo’s rotation, in which a ring particle completes one orbit for every three rotations of the central body [Salo2024].
 While effective for C1R, this resonance drives an outward migration of the ring, ultimately risking its departure from the resonance and subsequent dispersion.
-A small external shepherd moon has been hypothesised to prevent such drift and maintain C1R’s location [33].
-If present, this moonlet might share its orbit with C2R and act as a source of dust, analogous to the small satellites observed in Saturn’s and Uranus’s ring systems.
-This scenario is further supported by our estimate of C2R’s dynamical spreading timescale: based on JWST measurements, we find a median value of only 0.36 years (1 range: 0.23–0.77 years; see Methods, Section 1.6), implying that without confinement or replenishment, the ring should have dispersed since its discovery in 2013.
-A co-orbiting shepherd satellite could thus play a dual role—maintaining the ring’s sharp edges while continuously supplying fresh material [17].
-However, another recent dynamical study about the effect of solar radiation pressure on the stability of a small-body rings also shows that for highly tilted rings, such as around Chariklo, there are two plausible particle size range that can be stable: m and m, continuing to allow for both possible options [18].
-The emerging picture suggests that the two scenarios for C2R—rapid erosion and a dusty, micron-scale particle population—are not mutually exclusive.
-A ring composed of sub-micron grains may be inherently short-lived, as such particles are particularly susceptible to dynamical processes that promote dispersal.
-We therefore consider both effects jointly, as they may together account for the observed decline in opacity and signal strength of the C2R ring.
-If the outer ring is indeed degrading over a timescale of just a few years, this would mark it as a transient structure, potentially representative of a broader class of ephemeral rings among small bodies, as hinted by episodic activity and temporary bound structures around another Centaur, Chiron [10].
-If instead the C2R ring is linked to Chariklo’s activity as a Centaur-like object, we would again expect small grains, as cometary dust is predominantly sub-micron in size [34, 35].
-Crucially, a single occultation in visible light of a sufficiently bright star could distinguish between these scenarios: if the opacity matches values from 2013–2017, grain-size effects are favoured; but if the occultation yields significantly lower opacities or no detection at all, it would instead support the interpretation of an actively fading ring.
-Either outcome would raise fundamental questions about the longevity of dusty rings and the fate of metre-sized material subject to slow but persistent erosion over decadal timescales.
+A small external shepherd moon has been hypothesised to prevent such drift and maintain C1R’s location [Salo2024].
+If present, this moonlet might share its orbit with C2R and act as a source of ring material, analogous to the small satellites observed in Saturn’s and Uranus’s ring systems.
+This scenario is further supported by our estimate of C2R’s dynamical spreading timescale: based on JWST measurements, we find a median value of only 0.34 years (1 range: 0.21–0.96 years; see Methods, Section 1.6), implying that without confinement or replenishment, the ring should have dispersed since its discovery in 2013.
+A co-orbiting shepherd satellite could thus play a dual role—maintaining the ring’s sharp edges while continuously supplying fresh material [Sickafoose2024].
+However, another recent dynamical study about the effect of solar radiation pressure on the stability of a small-body rings shows that for highly tilted rings, such as the one around Chariklo, there are two plausible particle size ranges that can be stable: m and m [Regaly2025].
+If the outer ring is indeed degrading over a timescale of just a few years, this would mark it as a transient structure, potentially representative of a broader class of ephemeral rings among small bodies, as hinted by episodic activity and temporarily bound structures around another Centaur, Chiron [Ortiz2023].
+A single near-future occultation measurement in visible light (different from JWST wavelengths) could confirm the fading of C2R and the strengthening of C1R.
 Ultimately, our JWST observations unveil a new and unexpected level of complexity in minor body ring systems—one that challenges existing paradigms and opens an unprecedented window into the transient architectures sculpting the outer Solar System.
@@
 C2R was not detected in the F322W2 filter due to a low signal-to-noise ratio.
-1st and 2nd contact refer to before and after the closest approach, respectively.
+1st and 2nd contacts refer to the ones before and after the closest approach, respectively.
 Values are presented with uncertainties: asymmetric uncertainties are indicated with upper and lower limits (), while symmetric uncertainties are shown with a single value following the symbol.
@@
 1.1 Predicting and observing stellar occultations from space
-In June 2020, the first-ever predicted stellar occultation observation using a space telescope was obtained with the Characterising ExOPlanets Satellite (CHEOPS) [36].
-This observation demonstrated the feasibility of predicting and observing stellar occultations caused by distant solar system minor bodies using space telescopes: these facilities are unaffected by weather conditions and can extend the range of observations beyond the Earth’s surface.
-The use of space telescopes also facilitates the detection of material surrounding small bodies, as it enables photometric accuracy comparable to that achieved by ground-based telescopes with much larger apertures.
-We also have access to stellar occultations of fainter stars than from the ground, multiplying the number of possible events.
-JWST is orbiting close to the second Sun-Earth Lagrange point (L2) [37, 38].
-L2 is an unstable equilibrium point, meaning constant manoeuvres must be made to keep the telescope in orbit around it [39]: such manoeuvres are usually made 2–3 times a month.
+In June 2020, the first-ever predicted stellar occultation observation using a space telescope was obtained with the Characterising ExOPlanets Satellite (CHEOPS) [Morgado2022].
+This observation demonstrated the feasibility of predicting and observing stellar occultations by distant solar system minor bodies using space telescopes: these facilities are unaffected by weather and can extend the range of observations beyond Earth’s surface.
+The use of space telescopes also facilitates the detection of material surrounding small bodies, enabling photometric accuracy comparable to that achieved by ground-based telescopes with much larger apertures.
+We also have access to stellar occultations of fainter stars than we do from the ground, thereby multiplying the number of possible events.
+JWST is orbiting close to the second Sun-Earth Lagrange point (L2) [Gardner2023, McElwain2023].
+L2 is an unstable equilibrium point, meaning constant manoeuvres must be performed to keep the telescope in orbit around it [Menzel2023]: these are usually performed 2–3 times a month.
 Consequently, the predicted orbital path of JWST needs to be frequently updated (J.
 Giorgini, personal communication, June 22, 2022).
-For all the above reasons and due to the complex motion of the space telescope around L2, predicting stellar occultations by distant minor bodies visible from JWST is a very challenging task.
+For all the above reasons, and given the complex motion of the space telescope around L2, predicting stellar occultations by distant minor bodies visible from JWST is a very challenging task.
 Extended Data Figures shows the difference between two consecutive releases of the JWST ephemeris, obtained on September 27, 2022.
-In just under six months from the initial epoch, the difference between the two releases reached approximately 10,000 km, which is larger than the Earth’s radius.
+In just under six months from the initial epoch, the difference between the two releases reached approximately 10,000 km, exceeding the Earth’s radius.
 In the absence of formal positional uncertainty, the plot illustrates the trustworthiness of a stellar occultation prediction using JWST far from the initial epoch.
-To predict stellar occultations that JWST could potentially observe, we used the Stellar Occultation Reduction and Analysis (SORA) software package [40] combined with the Gaia DR3 catalogue [41, 42] and the NIMA ephemeris [43] for the solar system bodies.
+To predict stellar occultations that JWST could potentially observe, we used the Stellar Occultation Reduction and Analysis (SORA) software package [GomesJunior2022] combined with the Gaia DR3 catalogue [Gaia2016, Gaia2023] and the NIMA ephemeris [Desmars2015] for the solar system bodies.
 The strategy we developed to run the predictions is as follows.
 We selected the 40 largest TNOs/Centaurs (36 TNOs and 4 Centaurs), then ran predictions for these 40 targets every week for Gaia DR3 stars up to a G-magnitude of 20.
-This frequency of re-running the predictions was necessary due to the station-keeping manoeuvres applied to JWST, which changed the position (ephemeris) of the space telescope.
-Each week, we reviewed the updated predictions to identify favourable events, refined these predictions with new astrometric data obtained from the ground, and reran the predictions for those objects.
+This re-running of the predictions was necessary because of the station-keeping manoeuvres applied to JWST, which changed its position (ephemeris).
+Each week, we reviewed the updated predictions to identify favourable events, refined them with new ground-based astrometric data, and reran the predictions for those objects.
 Finally, when we were confident in the prediction, we activated the target of opportunity (ToO) to observe the occultation from JWST.
-Since this was a non-disruptive ToO, it had to be activated at least 14 days in advance of the event.
+Since this was a non-disruptive ToO, it had to be activated at least 14 days before the event.
 Using the described strategy, we identified the occultation by the ringed Centaur Chariklo on 18 October 2022 (UT), with a low relative velocity (2.5 km/s) and a star bright enough to be observed by JWST (K = 15.7 mag) as the most promising event.
 We first identified this occultation on 19 August 2022 (blue line in Extended Data Figures).
-Then we reran the prediction every week to account for changes in JWST’s orbit caused by station-keeping manoeuvres.
+Then we reran the prediction each week to account for changes in JWST’s orbit caused by station-keeping manoeuvres.
 Our most recent prediction was made on the day of the occultation, just a few hours before the event (green line in Extended Data Figures).
@@
 This observation was both scientifically valuable and technically demanding.
-It was conducted under JWST’s Solar System Guaranteed Time Observations (GTO) program #1271, ToO TNOs: Unveiling the Kuiper belt by stellar occultations (PI: Pablo Santos-Sanz) [44].
+It was conducted under JWST’s Solar System Guaranteed Time Observations (GTO) program #1271, ToO TNOs: Unveiling the Kuiper belt by stellar occultations (PI: Pablo Santos-Sanz) [SantosSanz2017].
 On 18 October 2022, between 12:36:04 and 13:50:27 UT, we observed the event with the JWST NIRCam instrument, targeting the star Gaia DR3 6873519665992128512 (RA 20h 21m 46s, Dec -16∘ 26′ 22′′; G = 17.5, K = 15.7).
@@
 As anticipated from the occultation geometry (Extended Data Figures), the main body of Chariklo did not intercept the line of sight, and no additional dips beyond the 2 noise level were detected, even when the complete data set was examined.
-Fluxes from Chariklo and the target star were also measured separately in both filters, using images acquired before the occultation when the two sources were spatially resolved (Extended Data Figures, middle Panel).
+Fluxes from Chariklo and the target star were also measured separately in both filters, using images acquired before the occultation when the two sources were spatially resolved (Extended Data Figures, left Panel).
 These measurements were used to calibrate the relative contributions of Chariklo and the star to the combined flux, a crucial step in deriving the ring parameters discussed in Section 1.4.
@@
 In the obtained light curves, the main, densest, and most opaque ring of Chariklo is clearly visible.
-This ring is called C1R, following the original notation [7].
-The C1R occultation light curve displays obvious diffraction spike features immediately before and after the ingress and egress, respectively.
-These spikes provide strong evidence that the ring has sharp edges and is tightly confined, similar to the rings of Uranus [23, 24].
-However, the dynamical mechanism responsible for this confinement is still unknown and is currently under discussion [16, 45, 46, 33].
+This ring is called C1R, following the original notation [Braga-Ribas2014].
+The C1R occultation light curve displays prominent diffraction spike features immediately before and after the ingress and egress, respectively.
+These spikes provide strong evidence that the ring has sharp edges and is tightly confined, similar to the rings of Uranus [Elliot1977, Elliot1987].
+However, the dynamical mechanism responsible for this confinement is still unknown and is currently under discussion [Sicardy2019, Sicardy2020a, GiuliattiWinter2023, Salo2024].
 The other ring (referred to as C2R) is surprisingly not clearly detected in the F150W2 or F322W2 light curves, as shown in Extended Data Figures.
 Since a ring partially occults the starlight, we fit an occultation model with a monolayer abrupt edge, including the effects of Fresnel diffraction, finite bandwidth, exposure time, and the Gaia DR3 stellar diameter.
-We use the procedures within the SORA package [40].
+We use the procedures within the SORA package [GomesJunior2022].
 From this analysis, we obtained the time when the star disappeared and reappeared behind Chariklo’s rings ( and ) and the observed apparent ring opacity (), which measures the fractional drop of stellar flux.
@@
 This projection was executed using the SORA package, and its result is shown in Figure 1.
-Assuming Chariklo’s parameters (body and rings) as published by [15], we fitted the centre position of Chariklo as viewed by JWST at the mid-point of the occultation (18 October 2022 at 13:37:42.18 UT), which resulted in
-That meant an appulse with the closest approach of 132.2 km, in contrast to Chariklo’s volume-equivalent radius of 124.8 km [15].
+Assuming Chariklo’s parameters (body and rings) as published by [Morgado2021], we fitted the centre position of Chariklo as viewed by JWST at the mid-point of the occultation (18 October 2022 at 13:37:42.18 UT), which resulted in
+That meant an appulse with the closest approach of 132.2 km, in contrast to Chariklo’s volume-equivalent radius of 124.8 km [Morgado2021].
 Initially, the parameters of both rings (C1R and C2R) were obtained in the sky plane simultaneously, just before and after the closest approach.
-We used the formalism presented by [14] to project these parameters to the ring plane so we could obtain the relevant parameters, such as the ring radial width (), the normal opacity (), and the normal optical depth ().
+We used the formalism presented by [Berard2017] to project these parameters to the ring plane so we could obtain the relevant parameters, such as the ring radial width (), the normal opacity (), and the normal optical depth ().
 Assuming a monolayer ring, the observed ring opacity normal to the ring () can be described as
-alternatively, if we assume a poly layer ring, is defined as [47]
+alternatively, if we assume a poly layer ring, is defined as [Elliot1984]
 where is the ring opening angle, and (resp.
@@
 edge-on) viewing.
-On the other hand, considering that the Airy scale is larger than the width of the ring as seen in the sky plane [48], a factor is needed to compute the fraction of lost light; in that case
+On the other hand, considering that the Airy scale is larger than the width of the ring as seen in the sky plane [Boissel2014], a factor is needed to compute the fraction of lost light; in that case
 for a monolayer or polylayer ring, respectively.
 The integrals of and over the radial width () of the ring define the equivalent width () and the equivalent depth () of the profile.
-These values ( and ) are proportional to the amount of material present in the profiles for each case [14].
+These values ( and ) are proportional to the amount of material present in the profiles for each case [Berard2017].
 The ring parameters are obtained in Table 1.
@@
 1.5 Sampling a ring with azimuthally variable structure
-In the case of a ring with an azimuthally variable structure, we may observe different normalized opacities () or equivalent with () values at the 1 and 2 contacts during an occultation event, as well as variations between multiple occultations.
+In the case of a ring with an azimuthally variable structure, we may observe different normal opacities () or equivalent with () values at the 1 and 2 contacts during an occultation event, as well as variations between multiple occultations.
 If the ring opacity is assumed to be wavelength-independent, the observed decrease/increase in opacity and/or in during the 2022 JWST events—relative to previous occultations—could be explained by the possibility that both ingress and egress sampled low/high-density segments of the ring.
@@
 Therefore, the probability of randomly sampling a pair of such low values in just one of eight events, as calculated from the simulation, represents an upper limit, with a probability of obtained in our simulation.
-In a similar way, we obtain a probability of p 0.004, assuming that we sample the ring close to the density maxima in the JWST measurements, and below these values in the pre-JWST cases (Extended Data Tables).
+Similarly, we obtain a probability of p 0.004, assuming that we sample the ring near the density maxima in the JWST measurements and below these values in the pre-JWST cases (Extended Data Tables).
 While azimuthal fragmentation is plausible—especially given the existence of fragmented rings in the trans-Neptunian region, e.g.
-in the case of Quaoar [36, 22]—these very low probabilities suggests that the observed JWST opacity drop (C2R) or increase (C1R) is unlikely to be solely due to this effect.
-The results have been found to be independent of the parameter for 3.
+in the case of Quaoar [Morgado2022, Proudfoot2025]—these very low probabilities suggest that the observed JWST opacity drop (C2R) or increase (C1R) is unlikely to be solely due to this effect.
+The results are independent of the parameter for 3.
 Here, we assumed that the azimuthal structures of the two rings are independent.
 1.6 Spreading timescale of Chariklo’s C2R Ring
-We compute the characteristic spreading timescale of Chariklo’s second ring (C2R) using the standard expression for a narrow, self-gravitating ring in resonance [49]:
+We compute the characteristic spreading timescale of Chariklo’s second ring (C2R) using the standard expression for a narrow, self-gravitating ring in resonance [DePater2018]:
 where is the gravitational constant, the mass of Chariklo, the ring semi–major axis, and its radial width. Chariklo’s mass is derived from its bulk density and ellipsoidal volume:
-We adopted ellipsoidal semi-axes from Morgado et al. (2021)[15]: km, km and km and a mean bulk density of .
+We adopted ellipsoidal semi-axes from Morgado et al. (2021)[Morgado2021]: km, km and km and a mean bulk density of .
 From our JWST occultation measurements of the C2R ring’s and contact radii ( and km), we obtain the weighted mean:
-To account for the uncertainty in radial width, we uniformly sampled over the KDE-inferred interval [0.40, 3.25] km.
-A Monte Carlo simulation with realisations yielded a median spreading timescale of s (0.36 yr), with a 1 confidence interval between s (0.23 yr) and s (0.77 yr).
+To account for the uncertainty in the radial width, we uniformly sampled over the interval defined by the minimum and maximum C2R widths measured across all occultations, [0.10, 3.72] km (see Extended Data Tables).
+A Monte Carlo simulation with realisations yields a median spreading timescale of s (0.34 yr), with a 1 confidence interval ranging from s (0.21 yr) to s (0.96 yr).
 1.7 Radiative Transfer Modeling
-After the striking marginal (at F150W2) and non-detection (at F322W2) of the C2R in the JWST filters, wavelength-dependent observables induced by micron- and submicron-sized grains became a real concern in explaining the measured discrepancies against previous occultations.
-Here, we perform a thorough analysis of the compatibility of dusty ring models for both C1R and C2R.
-Our main goals are to investigate which material in our models can replicate these observations and what we can say about the possible presence or absence of water ice in the rings.
-To explore whether a potential wavelength-dependent effect is responsible for the observed changes in the measured opacities, we conducted a radiative transfer analysis, assuming that the ring did not change during the observation period (2013-2022) and that all observational chords sampled the same amount of material, after correcting for the opening angle change due to Chariklo’s orbital motion around the Sun.
+To explore whether a potential wavelength-dependent effect is responsible for the observed changes in the measured opacities for C1R, we conducted a radiative transfer analysis, assuming that the ring did not change during the observation period (2013-2022) and that all observational chords sampled the same amount of material, after correcting for the opening angle change due to Chariklo’s orbital motion around the Sun.
 We used a slab-like, two-component model, where we mixed materials to match their common optical properties with the measured occultation normal opacities () or, equivalently, their normal optical depths ().
@@
 A single-component model is obtained if is chosen.
-To calculate the extinction coefficients, we used the OPTOOL package [50], which provides direct absorption and scattering coefficients ( and in units of cm2/g) for different types of materials and grain sizes.
-Previous studies based on reflectance spectra between 1998 and 2013 proposed an icy ring system around Chariklo and found that it may contain 20% water ice and 80% other materials, such as silicates, organics and small quantities of carbon [19], which motivated our choice of materials to investigate: water ice, carbon and silicates.
-OPTOOL offers both amorphous and crystalline options for most materials, and we found no significant differences between these options for our measurements.
+To calculate the extinction coefficients, we used the OPTOOL package [optool], which provides direct absorption and scattering coefficients ( and in units of cm2/g) for different types of materials and grain sizes.
+Previous studies based on reflectance spectra between 1998 and 2013 proposed an icy ring system around Chariklo and found that it may contain 20% water ice and 80% other materials, such as silicates, organics and small quantities of carbon [Duffard2014], which motivated our choice of materials to investigate: water ice, carbon and silicates.
+OPTOOL offers both amorphous and crystalline options for most materials, and we found no significant differences between them in our measurements.
 Therefore, we used the amorphous option in our analysis.
-We tested different kinds of silicates, such as pyroxene with various magnesium contents, but found minimal impact on the solutions.
+We tested different types of silicates, including pyroxene with varying magnesium content, but found minimal impact on the solutions.
 Therefore, we used a pyroxene model with 100% magnesium content, referred to as silicates hereafter.
-In general, although is a more observationally constrained value, compositional information is carried by the opacity estimates.
+In general, although is more observationally constrained, compositional information is carried by the opacity estimates.
 Model calculations (see Equation 9) can provide only optical depth values for direct comparison.
@@
 where is the wavelength-specific opacity, obtained from the model optical depths as = .
-First, we used single-grain models with sizes between 0.2 and 10 m.
+First, we used single-grain models with sizes ranging from 0.2 to 10 m.
 Sizes larger than 10 m do not show significant differences in extinction properties at different wavelengths in the visible to near-infrared range.
-We note that for planetary rings, we expect the grain size range to show a power-law distribution with the exponent around three [51].
-However, with the accuracy of the occultation measurements available, our primary goal was to identify the most dominant grain sizes.
+We note that for planetary rings, we expect the grain size range to show a power-law distribution with the exponent around three [Brilliantov2015].
+However, given the available accuracy of the occultation measurements, our primary goal was to identify the dominant grain sizes.
 In each case, we determined the best-fit models using least squares minimisation.
 The values were calculated as:
-where = with being the normal measured opacity, and is its uncertainty in the different filters.
+where = denotes the normal fractional transmission, with the measured normal opacity, and its associated uncertainty in the different filters.
 After identifying the best-fit grain sizes, we mixed the single-grain solutions using two approaches: (1) mixing the same material with two-grain sizes close to each other and (2) mixing two materials with their preferred grain sizes.
 The cases with the lowest values are considered as our compositional solutions for the rings.
-Even though there have been a large number of occultation measurements since 2013, only a few sites could capture the very narrow C2R ring.
-With their remarkable sensitivity and sampling abilities, the JWST measurements provided a unique opportunity to enrich the existing observations of the secondary ring C2R.
-Extended Data Tables shows all the available measurements from 2013 on the C2R ring, where the JWST ones are the most well-sampled of all those measurements.
-Some of the earlier C2R measurements contain only one or two data points, resulting in an unresolved detection of the ring.
-Therefore, it is essential to have a good understanding of the , and radial width values from previous observations and their associated uncertainties in the probability space.
-The most important of all is the VLT measurement in 2017 using the band, which samples the same, albeit smaller, piece of the JWST F150W2 filter range.
-Panel (a) in Extended Data Figures presents a comparative analysis of Chariklo’s ring system as observed with the VLT in 2017 (at 2.1 m, very close to the JWST F150W2 band) and with JWST in 2022.
-The radial brightness profiles reveal consistent locations and morphologies of the C1R and C2R ring components across both epochs and instruments, with only minor differences attributable to changes in viewing geometry and filter transmission curves.
-The JWST observations in the F150W2 and F322W2 filters provide higher signal-to-noise ratios, allowing finer sampling of the ring structure and sharper detection of the main features compared to ground-based data.
-Notably, however, while a stronger detection of C2R was expected with JWST, the opposite trend is observed: the C2R signal appears significantly weaker, especially in the F322W2 filter, where the z-score of the 1 and 2 contacts reaches only ~0.31 and ~0.23, respectively.
-Even in F150W2, where the JWST detection is somewhat clearer (~1.70 and ~1.73), it remains below the VLT levels (~4.42 and ~3.36).
-Nevertheless, the previously published opacity values for this observation appear asymmetrical at first glance.
-The 1 contact has a low opacity and well-constrained error margin (), which aligns well with the JWST measurements.
-However, the second contact appears completely different, with considerable uncertainty () -see Extended Data Tables.
-Figure Extended Data Figures Panel (b) shows the actual distributions of the possible apparent opacity values of this VLT occultation.
-Below the red lines, the 1- and 3- regions can be found.
-This is a typical result of fitting the occultation light curves using a Monte Carlo approach with different possible models and evaluating them.
-For the 1 contact, we can confirm that the value is fairly constrained.
-Still, for the second one, even though there is (numerically) a best-fitted value (), there is little to no difference between the values within a wide range (which resulted in error bars).
-This clearly demonstrates that, due to the low resolution of the VLT measurement, we cannot determine a single value for , but rather a range that, as with the 1 contact, overlaps with the JWST F150W2 region.
-Because the previously published individual opacity and width values may not accurately represent the most likely values within their given error bars, we used a probabilistic approach to examine the typical radial width values of the C2R ring.
-To identify the most representative radial widths () derived from the occultation light curve fits, we performed a kernel density estimation (KDE) using all valid C2R measurements.
-The resulting probability density function, shown in Figure Extended Data Figures, exhibits two distinct peaks that characterise the most probable width regimes ().
-These KDE-derived peaks guided the recalculation of the normal optical depth , using the relation:
-where is the equivalent width obtained from the light curve fitting.
-When the measured of an observation overlapped with a KDE peak, that value was directly used in Equation 12.
-In contrast, if the measured did not intersect any peak, a constrained search within its asymmetric uncertainty bounds was performed to select the value closest to the nearest KDE maximum, while ensuring it remained within the original asymmetric limits.
-From the set of candidate values yielding within the observational error limits, the extremal value (either minimum or maximum) was selected based on the direction of the nearest peak.
-This procedure was applied, for instance, to observations such as OPD (both contacts), Springbok (second contact), and Gifberg (second contact), which do not directly overlap with either KDE peak.
-This approach avoids arbitrary resampling and preserves both physical consistency and fidelity to observational uncertainties.
-By applying this method across all measurements, we derived new, more probable normal optical depth values, , constrained strictly within the original asymmetric uncertainties.
-The resulting distribution, shown in Figure 3, illustrates the original uncertainty ranges along with the resampled values (Extended Data Tables).
-The KDE-based adjustment significantly reduces the spread of the data points, highlighting consistent trends that were previously obscured.
-For instance, the two VLT observations, which originally showed substantial differences in (see Extended Data Tables), now exhibit closer agreement once KDE-informed widths are applied.
-However, the second, much larger VLT measurement also covers a wider potential opacity range with an overlap with the first, well-constrained value, which also agrees well with the JWST F150W2 observation.
-Our analyses indicate a reasonable likelihood that the two 2017 VLT measurements exhibit similar opacities to the 2022 JWST measurements, suggesting a silicate-like composition consisting of small grains.
-Considering also the Danish observations, the preferred silicate solutions for the C2R ring are between 0.2 – 0.5 m.
-The two Danish and the two JWST observations are crucial because, for each instrument, the observations in two filters were parallel in time and sampled the same part of the ring.
-Therefore, they are not sensitive to possible inhomogeneities in the ring.
-The bottom Panel of Figure 3 presents the comparison between the resampled (and original) observations and the models.
-The normal opacities transformed into relative intensities as , showing the fractional dimming of the occulted star, which now can be directly compared to the models.
-For the non-detection in the F322W2 filter, we indicated a violet range set to a relative intensity of 1.0, where the width of the region represents the uncertainty of the non-detection – larger signals should have been detected.
-We created pyroxene models with a magnesium content of 100% using 0.2, 0.3, 0.5, 1.0 and 3.0 m grain sizes, marked as red lines in the bottom panel of Figure 3.
-These silicate models can capture the general trend of the C2R opacities, which suggests that opacities in the visible wavelength are more prominent than the near-infrared ones.
-This decreasing signal strength aligns well with the fact that the smaller the grain size of the silicate models, the greater their observable opacity in the visible range.
-Therefore, we suggest a compatible dusty ring model for C2R dominated by silicate grains in the range of m.
-We also note that the optical properties of the pyroxene models are also suitable for explaining an asymmetric detection in the two JWST filters, as based on these models, we expect a higher signal in the F150W2 compared to F322W2.
-Since the JWST measurements in these filters were taken at the same time, sampling the same ring segment, such a silicate-dominated dusty ring may also contribute to an unexpected difference between the two filters.
-All these models show that in the F322W2 band, no significant signal can be detected, if grain sizes are kept below 3 m, as represented by the red curves confined to the violet-coloured region.
-The sky blue lines correspond to the water models with grain sizes of 0.5, 1.0 and 4.0 m, which display a distinctive signature feature at around 3 m.
-This particular feature is revisited in Figure 4, where we mainly focus on the JWST filter ranges.
-The middle panel of Figure 4 clearly shows that the smaller the grain size, the larger the feature itself.
-For grains larger than 2-5 in size, the curves flatten, distributing nearly equal contribution among all wavelengths.
-Fortunately, the wavelength range of this water signature is entirely covered by the F322W2 passband, offering a natural selection between the suggested models in this scenario.
-Taking the transmission profiles of each filter into account (top Panel in Figure 4), water ice grains of m would make a notable contribution in the F322W2 filter, relative to the F150W2 one.
-Pure water ice rings, similar to those around Saturn [30], cannot explain the measurements for either sub-micron or larger particles.
-Although our numerically best fit model for all the data is clearly the pure pyroxene one (see Extended Data Figures), the finite resolution of the measurements leaves room for other, similarly suitable mixture models.
-Therefore, the presence of a certain amount of water ice remains compatible with the measurements.
-However, in this purely wavelength-dependent scenario, the exact amount of water ice in C2R heavily depends on the chosen best-fit silicate model and the assumed water grain sizes.
-The bottom panel of Figure 4 exhibits silicate models with grain sizes of m, combined with different amounts of m, (solid) or m (dotted) ice grains.
-Besides the exact differences, what is the same in all models is that water content above 10-30% would lead to a decrease in the observed visual-infrared opacity discrepancy and a significant increase in the contribution of the F322W2 filter, which would then have been observed along with the marginal detection in F150W2.
-These estimates of water content are also very similar to those previously suggested for the C1R ring [19].
-To investigate the wavelength-dependent opacity of C1R, we performed a similar, simple radiative transfer calculation assuming a homogeneous layer for each occultation event, composed of a mixture of two materials with different compositions and grain sizes.
+Here, we perform a thorough analysis of the compatibility of dusty ring models for C1R, assuming a homogeneous layer for each occultation event, composed of a mixture of two materials with different compositions and grain sizes.
 The pre-JWST data, including the 2017 VLT/Ks measurement at , can be well fitted with a mixture of silicates (olivine or pyroxene) with grain sizes and water ice with grain sizes between 1–10 (blue curve in Figure 3).
@@
 However, grains larger than would produce nearly wavelength-independent opacity over the full range considered, regardless of composition, suggesting that the grains responsible for most of the extinction in C1R are likely smaller than this threshold.
-When the JWST data are included, however, no combination of two materials and grain sizes provides a satisfactory fit across all datasets—particularly the JWST F150W2, F322W2, and VLT/Ks bands.
+When JWST data are included, however, no combination of two materials and grain sizes provides a satisfactory fit across all datasets—particularly the JWST F150W2, F322W2, and VLT/Ks bands.
 The best-fit models in this case (red curve in Figure 3) require predominantly silicates with 1–10 grain sizes, but the overall fit quality remains poor.
 This strongly suggests that the observed opacity variations cannot be explained solely by wavelength dependence.
-For comparison, spectroscopy of the unresolved Chariklo system indicates a composition of 20% water ice mixed with darker materials, probably silicates [19].
+For comparison, spectroscopy of the unresolved Chariklo system indicates a composition of 20% water ice mixed with darker materials, probably silicates [Duffard2014].
 Extended Data Figures
 Extended Data Tables
-| Telescope/Instrument/Filter | Date |
-| Danish-1.54m/DFOSC/Za | 2013-06-03 | ||||||
-| VLT/Hawk-I/Ksa | 2014-02-16 | ||||||
-| SAAO/SHOCa | 2014-04-29 | ||||||
-| Springbok/Raptor/Emptya | 2014-04-29 | ||||||
-| Hakos/Raptor/Clearb | 2017-04-09 | - | |||||
-| WabiLodge/Raptor/Clearb | 2017-04-09 | ||||||
-| WeaverRL/Raptor/Clearb | 2017-04-09 | ||||||
-| Tivoli/Raptor/Clearb | 2017-06-22 | ||||||
-| WdhBackes/Raptor/Clearb | 2017-06-22 | ||||||
-| Onduruquea/Raptor/Clearb | 2017-06-22 | ||||||
-| Outeniqua/Raptor/Clearb | 2017-06-22 | ||||||
-| WdhMeza/ASI178MM/Clearb | 2017-06-22 | ||||||
-| Hakos/Raptor/Clearb | 2017-06-22 | ||||||
-| Danish-1.54m/Redb | 2017-07-23 | ||||||
-| VLT/Hawk-I/Ksb | 2017-07-23 | ||||||
-| SSO/Andor/g/I+zb | 2017-07-23 | ||||||
-| ElRodeo/MerlinKite/Clearb | 2017-07-23 | ||||||
-| ElSalvador/Raptor/Clearb | 2017-07-23 | ||||||
-| Danish-1.54m/Visualb | 2017-07-23 | ||||||
-| LaSilla1m/Raptor/Clearb | 2017-07-23 | ||||||
-| OPD/Andor/Clearb | 2017-07-23 | ||||||
-| TolarGrande/Raptor/Clearb | 2017-07-23 | ||||||
-| IncaDeOro/Raptor/Clearb | 2017-07-23 | ||||||
-| OPD/Andor/Clearb | 2017-08-24 | ||||||
-| Glenlee/Watec/Clearb | 2020-06-19 | ||||||
-| JWST/NIRCam/F150W2c | 2022-10-18 | ||||||
-| JWST/NIRCam/F322W2c | 2022-10-18 |
 | Telescope/Instrument/Filter | Date | Ring | |||
@@
 | JWST/NIRCam/F322W2c | 18/10/2022 | C1R |
-| Telescope/Instrument/Filter | Date | |||||||
-| (m) | (km) | (km) | (km) | (km) | ||||
-| Danish-1.54m/DFOSC/Za | 03/06/2013 | 1.054 | ||||||
-| Springbok/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
-| Gifberg/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
-| Danish-1.54m/Visualb | 23/07/2017 | 0.510 | ||||||
-| Danish-1.54m/Redb | 23/07/2017 | 0.770 | ||||||
-| OPD/Andor/Clearb | 23/07/2017 | 0.640 | ||||||
-| VLT/Hawk-I/K | 23/07/2017 | 2.145 | ||||||
-| JWST/NIRCam/F150W2c | 18/10/2022 | 1.655 | ||||||
-| JWST/NIRCam/F322W2c | 18/10/2022 | 3.225 | - | - | - | - |
-| Telescope/Instrument/Filter | Date | |||||||
-| (m) | (km) | (km) | (km) | (km) | ||||
-| Danish-1.54m/DFOSC/Za | 03/06/2013 | 1.054 | ||||||
-| Springbok/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
-| Gifberg/Raptor/Emptya | 29/04/2014 | 0.565 | ||||||
-| Danish-1.54m/Visualb | 23/07/2017 | 0.510 | ||||||
-| Danish-1.54m/Redb | 23/07/2017 | 0.770 | ||||||
-| OPD/Andor/Clearb | 23/07/2017 | 0.640 | ||||||
-| VLT/Hawk-I/Ksb | 23/07/2017 | 2.145 | ||||||
-| JWST/NIRCam/F150W2c | 18/10/2022 | 1.655 | ||||||
-| JWST/NIRCam/F322W2c | 18/10/2022 | 3.225 |
+References.: a. [Berard2017], b. [Morgado2021], c. This work.
+| Telescope/Instrument/Filter | Date |
+| Danish-1.54m/DFOSC/Za | 2013-06-03 | ||||||||||||
+| VLT/Hawk-I/Ksa | 2014-02-16 | ||||||||||||
+| SAAO/SHOCa | 2014-04-29 | ||||||||||||
+| Springbok/Raptor/Emptya | 2014-04-29 | ||||||||||||
+| Hakos/Raptor/Clearb | 2017-04-09 | – | – | ||||||||||
+| WabiLodge/Raptor/Clearb | 2017-04-09 | ||||||||||||
+| WeaverRL/Raptor/Clearb | 2017-04-09 | ||||||||||||
+| Tivoli/Raptor/Clearb | 2017-06-22 | ||||||||||||
+| WdhBackes/Raptor/Clearb | 2017-06-22 | ||||||||||||
+| Onduruquea/Raptor/Clearb | 2017-06-22 | ||||||||||||
+| Outeniqua/Raptor/Clearb | 2017-06-22 | ||||||||||||
+| WdhMeza/ASI178MM/Clearb | 2017-06-22 | ||||||||||||
+| Hakos/Raptor/Clearb | 2017-06-22 | ||||||||||||
+| Danish-1.54m/Redb | 2017-07-23 | ||||||||||||
+| VLT/Hawk-I/Ksb | 2017-07-23 | ||||||||||||
+| SSO/Andor/g/I+zb | 2017-07-23 | ||||||||||||
+| ElRodeo/MerlinKite/Clearb | 2017-07-23 | ||||||||||||
+| ElSalvador/Raptor/Clearb | 2017-07-23 | ||||||||||||
+| Danish-1.54m/Visualb | 2017-07-23 | ||||||||||||
+| LaSilla1m/Raptor/Clearb | 2017-07-23 | ||||||||||||
+| OPD/Andor/Clearb | 2017-07-23 | ||||||||||||
+| TolarGrande/Raptor/Clearb | 2017-07-23 | ||||||||||||
+| IncaDeOro/Raptor/Clearb | 2017-07-23 | ||||||||||||
+| OPD/Andor/Clearb | 2017-08-24 | ||||||||||||
+| Glenlee/Watec/Clearb | 2020-06-19 | ||||||||||||
+| JWST/NIRCam/F150W2c | 2022-10-18 | ||||||||||||
+| JWST/NIRCam/F322W2c | 2022-10-18 |
+References.: a. [Berard2017], b. [Morgado2021], c. This work.
+| Telescope/Instrument/Filter | Date |
+| Danish-1.54m/DFOSC/Za | 03/06/2013 | ||||||||||||
+| Springbok/Raptor/Emptya | 29/04/2014 | ||||||||||||
+| Gifberg/Raptor/Emptya | 29/04/2014 | ||||||||||||
+| Danish-1.54m/Visualb | 23/07/2017 | ||||||||||||
+| Danish-1.54m/Redb | 23/07/2017 | ||||||||||||
+| OPD/Andor/Clearb | 23/07/2017 | ||||||||||||
+| VLT/Hawk-I/K | 23/07/2017 | ||||||||||||
+| JWST/NIRCam/F150W2c | 18/10/2022 |
+References.: a. [Berard2017], b. [Morgado2021], c. This work.
 Acknowledgements
@@
 These observations are associated with program #1271.
+All the JWST data used in this paper can be found in MAST: https://doi.org/10.17909/qaaf-8q45.
 This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium).
@@
 HBH and SNM also acknowledge support from NASA JWST Interdisciplinary Scientist grant 21-SMDSS21-0013.
+CK was supported by the K-138962 and TKP2021-NKTA-64 grants of the National Research, Development and Innovation Office (NKFIH), Hungary.
+We sincerely thank both anonymous referees for their exceptional and insightful contributions, which significantly improved the clarity and quality of this manuscript.